OpenStudy (anonymous):

Let F'(x) = sqrt (x+1) and F(0) = 3. Then

OpenStudy (anonymous):

Integrate it and substitute the values

OpenStudy (anonymous):

Remember to + c

OpenStudy (anonymous):

ok i'm kinda in a crunch mode for time... so could you show me?? I usually want hints only and work on it on my own.. but not for this question!

OpenStudy (anonymous):

ur saying I should integrate the first equation right?

OpenStudy (psymon):

\[\int\limits_{}^{}x ^{n}dx=\frac{ x ^{n+1} }{ n+1 }du\]

OpenStudy (anonymous):

Example f'(x) = -1 / x^2 f(1) = 10 f(x) = 1 / x + c f(1) = 1 / 1 + c = 1 + c = 10 c = 9 f(x) = 1 / x + 9

OpenStudy (psymon):

\[\int\limits_{}^{}(x+1)^{\frac{ 1 }{ 2 }}dx\]

OpenStudy (psymon):

Integrate that :P You get your answer, add + C, then you have the inital condition of f(0) = 3, which will allow you to solve for C and rewrite the full answer.

OpenStudy (anonymous):

hmmm ok... is it F(x) = 2/3(x+1)^(3/2)+7/3 .

OpenStudy (psymon):

Awesome job : )

OpenStudy (anonymous):

thanks!!!

OpenStudy (psymon):

^_^